OpenStudy (anonymous):

x^2-38x+326=0 please factor

OpenStudy (anonymous):

\[x=19\pm \sqrt{35}\]

OpenStudy (anonymous):

Just apply the quadratic formula

OpenStudy (anonymous):

use the quadratic formula \[x=\frac{-b \pm \sqrt {b^2-4ac}}{2a}, \] which A is the coefficient for the x^2 term B is the coefficient for the x term C is the coefficient for the x^0 term

OpenStudy (anonymous):

Yup, robtotey has the correct answer and Kynosis gives the formula

OpenStudy (anonymous):

Well it looks like I will need to start this from the beginning. This is new to me so I am lost. I have a word problem. It is about a ladder leaning against a bldg. distance from bottom of ladder to bldg is 18' less than the length of the ladder. How high up the side of the bldg is the top of the ladder if that distance is 1' less than the length of the ladder

OpenStudy (anonymous):

Maybe I set it up wrong. I started with \[(x-18)^2+(x-1)^2=x^2\]

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

I agree with you Geminiman

OpenStudy (anonymous):

finding the roots give you the factor of the equation indirectly. because equation are in the quadratic form is (x - n) ( x- n) with n being the root you have a factored equation. of \[( x - 19+√(35)) ( x - 19-√(35))\]

OpenStudy (anonymous):

ok. then I did this formula: \[x^2-36+324+x^2-2x+2=x^2\]

OpenStudy (anonymous):

I get x^2 - 38x + 325 = 0

OpenStudy (anonymous):

ok I got x^2338x+326=0

OpenStudy (anonymous):

Then you get (x - 25)(x - 13) = 0 which is x = 25, 13

OpenStudy (anonymous):

sorry typed it wrong. x^2-38x+326=0

OpenStudy (anonymous):

When you took (x - 1)^2 Gem you had x^2 - 2x + 2 and it is actually x^2 - 2x + 1

OpenStudy (anonymous):

ah ok that is why I got 326 instead of 325.

OpenStudy (anonymous):

Thank you all for your help